if the mean and the standard deviation of a continuous random variable that is normally distributed are 26 and 6, find an interval that contains 95% of the distribution
1
Expert's answer
2020-12-15T19:15:34-0500

According to the empirical rule:

95% fall within two standard deviations.

"= 26 ± 2 \times 6 \\

= 26 ± 12 \\

= (26-12, 26+12) \\

= (14, 38)"